Three Step Tariff. Determine the Total Tariff.
D: P = 400 - Q
SD: P = 40 + Q
Sw: P = 80
Tariff: P = 40
1
Expert's answer
2016-05-27T12:44:03-0400
Determine the SWL.
D: P = 400 - Q
SD: P = 40 + Q.
In equilibrium Qsd = Qd, so:
400 - Q = 40 + Q,
2Q = 360,
Q = 180 units.
P = 400 - Q = 400 - 180 = $220.
If the world supply (Sw) is: P = $80 and the tariff of P = $40 is imposed, then the new price is P = 80 + 40 = $120.
The new quantity demanded is:
120 = 400 - Q,
Q = 280 units.
The quantity supplied domestically is:
120 = 40 + Q,
Q = 80 units, so the amount of import is:
SWL = 280 - 80 = 200 units.